# ChuLo: Chunk-Level Key Information Representation for Long Document Processing

### v3

@@ Abstract
Transformer-based models have achieved remarkable success in various Natural Language Processing (NLP) tasks, yet their ability to handle long documents is constrained by computational limitations. Traditional approaches, such as truncating inputs, sparse self-attention, and chunking, attempt to mitigate these issues, but they often lead to information loss and hinder the model’s ability to capture long-range dependencies. In this paper, we introduce ChuLo, a novel chunk representation method for long document classification that addresses these limitations. Our ChuLo groups input tokens using unsupervised keyphrase extraction, emphasizing semantically important keyphrase based chunk to retain core document content while reducing input length. This approach minimizes information loss and improves the efficiency of Transformer-based models. Preserving all tokens in long document understanding, especially token classification tasks, is especially important to ensure that fine-grained annotations, which depend on the entire sequence context, are not lost. We evaluate our method on multiple long document classification tasks and long document token classification tasks, demonstrating its effectiveness through comprehensive qualitative and quantitative analyses.
@@ 1 Introduction
Transformer-based models (Vaswani et al., 2017), including LLMs (Radford, 2018; Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a; b; Chowdhery et al., 2023; Anil et al., 2023; Dubey et al., 2024), have achieved remarkable success across a wide range of Natural Language Processing (NLP) tasks, including Machine Translation, Text Summarization, Text Generation, and Text Classification. A key factor behind their success is the self-attention mechanism, which allows the model to capture long-range dependencies by computing the similarity between any two tokens and aggregating information accordingly. However, this mechanism incurs a quadratic computational cost in terms of both time and space, relative to input length. This computational burden makes it difficult for Transformer-based models to scale to long documents, limiting their application to real-world data with unrestricted document lengths.
To address this challenge, several approaches have been proposed for applying Transformer-based models to long documents while managing computational resources. One of them is truncating, where the model discards content exceeding a predefined input length. For instance, BERT (Kenton & Toutanova, 2019) processes up to 512 tokens, and LLaMa (Touvron et al., 2023a) handles up to 2048 tokens, with any additional content being ignored. Another one is sparse self-attention, which reduces computational complexity by restricting each query token to attend only to a subset of key tokens (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020; Wei et al., 2021; Li et al., 2023a). Lastly, chunking divides long documents into smaller, manageable segments that are processed independently by the model (Zhao et al., 2021; Zhang et al., 2022).
@@
The key contributions of this paper are as follows: 1) Novel Chunk Representation Method: We introduce ChuLo, a chunk representation method for long document understanding that leverages unsupervised keyphrase extraction to prioritize semantically important information, effectively reducing input length while preserving core content. 2) Enhanced Document and Token Classification: Our proposed method is designed to handle both document-level and token-level tasks, addressing the limitations of existing models in retaining fine-grained annotations and global context in long documents. 3) Scalable and Efficient Solution: ChuLo offers a scalable and efficient approach for long document processing, making it suitable for various NLP applications where handling long-range dependencies and context preservation are critical.
@@ 2.1 Long Document Understanding
Document understanding includes two directions: global understanding (e.g., document classification) and token-level understanding (e.g., named entity recognition). With Transformer-based models, document length impacts classification performance. Approaches, shown in Appendix A.1, to applying Transformer-based models on long document classification can be divided into input processing and Transformer architecture optimization. Input processing involves truncating and chunking. Truncating drops tokens exceeding the model’s input length (Park et al., 2022), while chunking segments documents into smaller parts processed separately. For example, Hierarchical Transformer (Pappagari et al., 2019) splits documents into non-overlapping chunks and computes chunk representations. RoR (Zhao et al., 2021) generates regional answers from chunks, which are combined for the final answer. However, neither considers the entire document context when chunking. Transformer architecture optimization includes two strategies: improving self-attention efficiency and incorporating RNN concepts. Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), and others (Roy et al., 2021) use sparse attention, restricting queries to focus on specific keys. Other methods (Peng et al., 2021; Wang et al., 2020; Choromanski et al., 2020) approximate self-attention with reduced complexity. Meanwhile, (Dai et al., 2019; Hutchins et al., 2022; Li et al., 2023b) integrate cache memory to update history information. These approaches involve performance-efficiency trade-offs, making it valuable to explore improving performance by reducing input length. In document NER, text length has received less attention. Recent research focuses on low-resource languages (Çetindağ et al., 2023; Mengliev et al., 2024), complex domain-specific texts (Park et al., 2023; Bhattacharya et al., 2023), prompt-based large model methods (Wang et al., 2023; Dagdelen et al., 2024; Hu et al., 2024), and multimodal NER (Yu et al., 2023; Zhang et al., 2023). Our work addresses these existing challenges by introducing a novel chunk representation that reduces input length while retaining key information, improving both classification and token-level tasks through semantic grouping of important phrases. Our method will preserve global and local features, achieving better efficiency and performance compared to existing models.
@@ 2.2 Unsupervised Keyphrase Extraction
Unsupervised keyphrase extraction automatically identifies representative phrases from a document to summarize its content without requiring labeled data (Hasan & Ng, 2014). Based on the features used, unsupervised methods can be categorized into three types: statistics-based, graph-based, and embedding-based (Kong et al., 2023). Statistics-based methods rank candidate phrases using features like TfIdf (El-Beltagy & Rafea, 2009), co-occurrence (Liu et al., 2009), and context statistics (Campos et al., 2020; Won et al., 2019). Graph-based methods construct a graph where nodes are candidate phrases and edges represent their relationships, such as TextRank (Mihalcea & Tarau, 2004) and its variants (Wan & Xiao, 2008; Bougouin et al., 2013; Florescu & Caragea, 2017; Yu & Ng, 2018). Embedding-based methods leverage distributional representations of words or sentences to achieve state-of-the-art performance, as seen in EmbedRank (Bennani-Smires et al., 2018), SIFRank (Sun et al., 2020), and PromptRank (Kong et al., 2023). Although these methods have shown effectiveness in capturing keyphrases, they focus on optimizing phrase extraction and ranking independently, rather than enhancing downstream tasks like long document representation. Our work bridges this gap by integrating unsupervised keyphrase extraction with chunk representation to improve long document understanding. We highlight the limitations of existing methods, which fail to preserve key content when reducing document length, and propose a novel approach that groups semantically important phrases to maintain critical information for downstream tasks.
@@ 3 ChuLo
We introduce a novel chunk representation method, ChuLo, tailored to enhance Transformer-based long document classification by effectively reducing input length while preserving semantic content. Our approach addresses the challenges posed by existing techniques like truncation and standard chunking, which often result in information loss and disrupted contextual dependencies. Specifically, we first segment the document into non-overlapping, fixed-length chunks to manage long input sequences. Next, we employ unsupervised keyphrase extraction to capture the semantically critical information across the document. By integrating these keyphrases into the chunk representations, we strategically assign higher weights to the keyphrase tokens, emphasizing essential content in each chunk. The resulting chunk representations are then used to train a Transformer-based chunk attention module, which ensures that the model retains the global context while focusing on important local information. This methodology not only mitigates the issue of information loss but also enables the model to handle long documents with improved efficiency and performance. The details of each component are in the following subsections, and the overall framework is in Figure 1.
@@ 3.1 Document Input Chunking
To effectively manage long document inputs, we employ a chunking strategy that reduces input length while preserving all relevant information. Transformer models, despite their success in capturing long-range dependencies through self-attention, suffer from quadratic computational complexity as input length increases (Ivgi et al., 2023; Beltagy et al., 2020). This complexity imposes limitations on the maximum input length and, consequently, on the amount of text the model can process at once. Common solutions such as truncating and sparse attention either disregard parts of the document (Lewis et al., 2020; Park et al., 2022) or limit the receptive field of individual tokens (Beltagy et al., 2020; Zaheer et al., 2020; Brown et al., 2020), leading to information loss. Our approach mitigates these issues by segmenting the document into non-overlapping chunks before feeding them into the model. This strategy enables complete self-attention among chunk, ensuring that all information is retained and enabling the model to process larger portions of the document context. Specifically, we first tokenize the document $D=(t_{0},t_{1},\dots,t_{l_{D}-1})$ and divide it into fixed-length chunks ${\mathbb{C}}_{D}=(C_{0},C_{1},\dots,C_{m-1})$, where $l_{D}$ is the number of the tokens, each chunk $C$ consists of $n$ tokens, and $m=\lceil{\frac{l_{D}}{n}}\rceil$ is the number of chunks. The incomplete chunks will be padded with the [PAD] tokens. The chunk size $n$ is a hyper-parameter controlling the degree of input length reduction. By grouping tokens this way, we maintain the integrity of the input content while alleviating the computational burden associated with processing long sequences.
@@ 3.2 Semantic Key Information Extraction
The fundamental reason for extracting keyphrases from the chunks, as defined in the document chunking step, is to maintain the integrity of the document’s semantic content while reducing input length. During chunking, the document is divided into smaller segments, which can inadvertently distribute important semantic information unevenly across chunks or even cause it to be diluted. Simply treating each chunk equally may lead to overlooking critical context essential for accurate document classification and token-level understanding. Identifying and highlighting critical phrases within these chunks ensures that the most relevant information is preserved and emphasized, allowing the model to focus on the core content even within a limited input space. This compensates for the information fragmentation caused by chunking and guides the Transformer’s attention mechanism to prioritize the most informative parts of the text, enhancing the model’s ability to capture the document’s overall meaning and relationships. Thus, extracting keyphrases from chunks is crucial for bridging the gap between document segmentation and semantic coherence, ultimately improving the effectiveness of the chunk-based representation for long document understanding. To achieve this, we extract semantically important keyphrases to identify the core content of the entire document. Since document understanding, such as document classification or token classification, inherently involves semantic understanding, it is crucial to highlight the most informative parts of the text to create meaningful chunk representations. By making the extracted keyphrases more salient, we can effectively emphasize the content that contributes most to the document’s overall meaning. Hence, we employ unsupervised keyphrase extraction methods, ensuring our approach remains adaptable across diverse domains without requiring annotated data. Building on the principles of PromptRank (Kong et al., 2023), we adapt and enhance its template-based approach to prioritize keyphrases that are contextually significant across the entire document. Our modified strategy, the Semantic Keyphrase Prioritization (SKP) Algorithm, leverages prompts to assess the importance of each candidate keyphrase, ensuring that semantically crucial information is highlighted for downstream document understanding. The details of this process are provided in Algorithm 1:
Algorithm 1 Semantic Keyphrase Prioritization (SKP) Algorithm
[…]
Parameter: Experimentally determined $\alpha$, $\gamma$
Output: Sorted keyphrases set ${\mathbb{K}}_{s}$
1:  Let ${\mathbb{S}}=\emptyset$, ${\mathbb{K}}_{s}=\emptyset$, $i=0$, $j=0$.
2:  Get the candidate phrases set: ${\mathbb{K}}=\mathcal{F}_{REG}(\mathcal{F}_{POS}(D))=\{k_{0},k_{1},\dots,k_{n-1}\}$
3:  Split $D$ into segments ${\mathbb{S}}=\{D_{0},D_{1},\dots,D_{m-1}\}$ to meet the input requirement of $\mathcal{F}_{E}$
4:  for  $i<n$ do
7:     for  $j<m$ do
9:     end for
11:  end for
12:  return ${\mathbb{K}}_{s}=Sort({\mathbb{K}},s)$
While PromptRank uses prompts to rank keyphrases across the first segment of the document determined by its encoder model, our SKP applies this concept at the entire document level to ensure that each chunk can preserve the most informative content for the entire document. After obtaining the sorted phrases set ${\mathbb{K}}_{s}$, we select top-n phrases as the keyphrases of the document, which can be regarded as ranked phrases according to their contextual significance within the entire document.
This approach emphasizes keyphrases during chunk representation, making critical semantic information more salient. Consequently, our method bridges the gap between document segmentation and semantic coherence by ensuring that key content is preserved and highlighted within the entire document, despite input length constraints. By integrating these keyphrase extraction techniques, our framework effectively identifies and emphasizes the most informative parts of a long document, resulting in improved chunk-based representations. This ensures that the Transformer retains critical context and relationships, ultimately enhancing its performance in long document classification and token-level understanding tasks.
3.3 Chunk Representation with Semantic Emphasis
@@
After extracting the semantically significant keyphrases, we construct a chunk representation that preserves and highlights this key information, ensuring that the chunk retains the core semantic content of the document. While chunking helps reduce the input length, it may also result in an uneven distribution of meaningful content across chunks. Thus, it is crucial to re-emphasize the importance of these keyphrases within the chunk to maintain semantic integrity. Our approach dynamically adjusts the representation of each chunk by assigning greater importance to keyphrase tokens, enabling the model to focus on the most relevant content during downstream processing.
To achieve this, we label the tokens corresponding to the extracted keyphrases in the original text as keyphrase tokens $T_{k}$, while other tokens are labeled as non-keyphrase tokens $T_{nk}$. Then, we feed these chunked tokens $t$ into the embedding layer to obtain their embeddings. The chunk embedding ${\bm{e}}_{C}$ is then computed using a weighted average of these token embeddings, as defined in Formula 1:
$\begin{cases}w_{t}=\begin{cases}a,\text{t is $T_{k}$}\\ b,\text{t is $T_{nk}$}\end{cases}\\ {\bm{e}}_{c}=\frac{\sum{w_{t}*{\bm{e}}_{t}}}{\sum{w_{t}}}\end{cases}$ (1)
Here, $w_{t}$ represents the weight assigned to each token $t$ in the chunk, where $a$ and $b$ are hyperparameters with $a>b.$ ${\bm{e}}_{t}$ denotes the embedding of token $t$, and ${\bm{e}}_{c}$ is the resulting chunk embedding that captures the weighted importance of keyphrase and non-keyphrase tokens. By assigning a higher weight $a$ to keyphrase tokens, we ensure that the resulting chunk representation emphasizes the most critical information while maintaining a compact input length.
@@
Finally, the chunk embeddings are fed into the Transformer-based model, allowing it to effectively leverage the enhanced chunk representations during long document classification or token-level understanding tasks. This method not only preserves the semantic coherence of the document but also allows the model to retain meaningful context and relationships, ultimately enhancing its performance on long document tasks.

### v4

# ChuLo: Chunk-Level Key Information Representation for Long Document Understanding

@@ Abstract
Transformer-based models have achieved remarkable success in various Natural Language Processing (NLP) tasks, yet their ability to handle long documents is constrained by computational limitations. Traditional approaches, such as truncating inputs, sparse self-attention, and chunking, attempt to mitigate these issues, but they often lead to information loss and hinder the model’s ability to capture long-range dependencies. In this paper, we introduce ChuLo, a novel chunk representation method for long document understanding that addresses these limitations. Our ChuLo groups input tokens using unsupervised keyphrase extraction, emphasizing semantically important keyphrase based chunks to retain core document content while reducing input length. This approach minimizes information loss and improves the efficiency of Transformer-based models. Preserving all tokens in long document understanding, especially token classification tasks, is important to ensure that fine-grained annotations, which depend on the entire sequence context, are not lost. We evaluate our method on multiple long document classification tasks and long document token classification tasks, demonstrating its effectiveness through comprehensive qualitative and quantitative analysis. Our implementation is open-sourced on GitHub [^1] https://github.com/adlnlp/Chulo.
ChuLo: Chunk-Level Key Information Representation
for Long Document Understanding
Yan Li1, Soyeon Caren Han2thanks: Corresponding Author, Yue Dai3, Feiqi Cao1, 1The University of Sydney, 2The University of Melbourne, 3The University of Western Australia 1{yali3816, fcao0492}@uni.sydney.edu.au, 2caren.han@unimelb.edu.au, 3yue.dai@research.uwa.edu.au
@@ 1 Introduction
Transformer-based models (Vaswani et al., 2017), including LLMs (Radford, 2018; Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a, b; Chowdhery et al., 2023; Anil et al., 2023; Dubey et al., 2024), have achieved remarkable success across a wide range of Natural Language Processing (NLP) tasks, including Machine Translation, Text Summarization, Text Generation, and Text Classification. A key factor behind their success is the self-attention mechanism, which allows the model to capture long-range dependencies by computing the similarity between any two tokens and aggregating information accordingly. However, this mechanism incurs a quadratic computational cost in terms of both time and space, relative to input length. This computational burden makes it difficult for Transformer-based models to scale to long documents, limiting their application to real-world data with unrestricted document lengths. To address this challenge, several approaches have been proposed for applying Transformer-based models to long documents while managing computational resources. One of them is truncating, where the model discards content exceeding a predefined input length. For instance, BERT (Kenton and Toutanova, 2019) processes up to 512 tokens, and LLaMa (Touvron et al., 2023a) handles up to 2048 tokens, with any additional content being ignored. Another one is sparse self-attention, which reduces computational complexity by restricting each query token to attend only to a subset of key tokens (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020; Wei et al., 2021; Li et al., 2023a). Lastly, chunking divides long documents into smaller, manageable segments that are processed independently by the model (Zhao et al., 2021; Zhang et al., 2022).
@@
The key contributions of this paper are as follows: 1) Novel Chunk Representation Method: We introduce ChuLo, a chunk representation method for long document understanding that leverages unsupervised keyphrase extraction to prioritize semantically important information, effectively reducing input length while preserving core content. 2) Enhanced Document and Token Classification: Our proposed method is designed to handle both document-level and token-level tasks, addressing the limitations of existing models in retaining fine-grained annotations and global context in long documents. 3) Scalable and Efficient Solution: ChuLo offers a scalable and efficient approach for long document understanding, making it suitable for various NLP tasks where handling long-range dependencies and context preservation is critical.
@@ 2.1 Long Document Understanding
Document understanding involves global understanding (e.g., classification) and token-level tasks (e.g., named entity recognition). Transformer-based models face performance issues with long inputs, addressed through input processing and architecture optimization. Input processing methods include truncating tokens beyond the input limit (Park et al., 2022) and chunking, as seen in Hierarchical Transformer (Pappagari et al., 2019) and RoR (Zhao et al., 2021), though these often neglect full document context. Architecture optimizations improve efficiency using sparse attention (Beltagy et al., 2020; Zaheer et al., 2020; Roy et al., 2021) or approximations (Peng et al., 2021; Wang et al., 2020; Choromanski et al., 2020). Other approaches incorporate RNN concepts, such as cache memory (Dai et al., 2019; Hutchins et al., 2022; Li et al., 2023b). These methods balance performance and efficiency, highlighting the need to reduce input length effectively.
For document NER, text length is less studied, with recent work addressing low-resource languages (Çetindağ et al., 2023; Mengliev et al., 2024), complex domains (Park et al., 2023; Bhattacharya et al., 2023), prompt-based methods (Wang et al., 2023; Dagdelen et al., 2024; Hu et al., 2024), and multimodal data (Yu et al., 2023; Zhang et al., 2023; Li et al., 2025). Our work tackles these challenges [^2] The summary of Long Document Understanding related works can be found in Appendix A.1 with a novel chunk representation that preserves semantic information while reducing input length, enhancing both classification and token-level tasks.
@@ 2.2 Unsupervised Keyphrase Extraction
Unsupervised keyphrase extraction identifies representative phrases to summarize content without labelled data (Hasan and Ng, 2014). Methods include statistics-based (e.g., TfIdf (El-Beltagy and Rafea, 2009), co-occurrence (Liu et al., 2009), and context statistics (Campos et al., 2020; Won et al., 2019)), graph-based (e.g., TextRank (Mihalcea and Tarau, 2004) and its variants (Wan and Xiao, 2008; Bougouin et al., 2013; Florescu and Caragea, 2017; Yu and Ng, 2018)), and embedding-based approaches (e.g., EmbedRank (Bennani-Smires et al., 2018), SIFRank (Sun et al., 2020), and PromptRank (Kong et al., 2023)). While effective, these methods prioritize phrase extraction and ranking over improving downstream tasks. Our work integrates keyphrase extraction with chunk representation to enhance long document understanding.
@@ 3 ChuLo
We propose ChuLo, a novel chunk representation method that enhances long document understanding by reducing input length while preserving semantic content. Unlike existing approaches like truncation and standard chunking, ChuLo minimizes information loss and maintains contextual dependencies. The method segments documents into non-overlapping chunks, integrates key semantic information using unsupervised keyphrase extraction, and assigns higher weights to keyphrase tokens in chunk representations. These enriched chunks train a Transformer-based chunk attention module, enabling efficient processing of long documents while retaining global and local context. Further details are provided in subsequent subsections, with the framework illustrated in Figure 1.
@@ 3.1 Document Input Chunking
To effectively manage long document inputs, we employ a chunking strategy that reduces input length while preserving all relevant information. Common approaches to long document processing, such as truncation and sparse attention, either disregard parts of the document(Lewis et al., 2020; Park et al., 2022) or restrict the receptive field of individual tokens (Beltagy et al., 2020; Zaheer et al., 2020; Brown et al., 2020), resulting in potential information loss. Our approach mitigates these issues by segmenting the document into non-overlapping chunks before feeding them into the model. It enables complete self-attention among chunks, ensuring that all information is retained and enabling the model to process larger portions of the document context. Specifically, we first tokenize the document $D=(t_{0},t_{1},\dots,t_{l_{D}-1})$ and divide it into fixed-length chunks $C_{D}=(C_{0},C_{1},\dots,C_{m-1})$, where $l_{D}$ is the number of the tokens, each chunk $C$ consists of $n$ tokens, and $m=\lceil{\frac{l_{D}}{n}}\rceil$ is the number of chunks. The incomplete chunks will be padded with the [PAD] tokens. The chunk size $n$ is a hyper-parameter controlling the degree of input length reduction. By grouping tokens this way, we maintain the integrity of the input content while alleviating the computational burden associated with processing long sequences.
@@ 3.2 Semantic Key Information Extraction
The fundamental reason for extracting keyphrases from the chunks, as defined in the document chunking step, is to maintain the integrity of the document’s semantic content while reducing input length. During chunking, the document is divided into smaller segments, which can inadvertently distribute important semantic information unevenly across chunks or even cause it to be diluted. Simply treating each chunk equally may lead to overlooking critical context essential for accurate document classification and token-level understanding. Identifying and highlighting critical phrases within these chunks ensures that the most relevant information is preserved and emphasized, allowing the model to focus on the core content even within a limited input space. This compensates for the information fragmentation caused by chunking and guides the Transformer’s attention mechanism to prioritize the most informative parts of the text, enhancing the model’s ability to capture the document’s overall meaning and relationships. Thus, extracting keyphrases from chunks is crucial for bridging the gap between document segmentation and semantic coherence, ultimately improving the effectiveness of the chunk-based representation for long document understanding. To achieve this, we extract semantically important keyphrases to identify the core content of the entire document. Since document understanding, such as document classification or token classification, inherently involves semantic understanding, it is crucial to highlight the most informative parts of the text to create meaningful chunk representations. By making the extracted keyphrases more salient, we can effectively emphasize the content that contributes most to the document’s overall meaning. Hence, we employ unsupervised keyphrase extraction methods, ensuring our approach remains adaptable across diverse domains without requiring annotated data. Building on the principles of PromptRank (Kong et al., 2023), we adapt and enhance its template-based approach to prioritize keyphrases that are contextually significant across the entire document. Our modified strategy, the Semantic Keyphrase Prioritization (SKP) Algorithm, leverages prompts to assess the importance of each candidate keyphrase, ensuring that semantically crucial information is highlighted for downstream document understanding. The details of this process are provided in Appendix A.2. In this way, our method emphasizes keyphrases during chunk representation, making critical semantic information more salient. Consequently, our method bridges the gap between document segmentation and semantic coherence by ensuring that key content is preserved and highlighted within the entire document, despite input length constraints.
3.3 Chunk Representation Production
@@
After extracting the semantically significant keyphrases, we construct a chunk representation that preserves and highlights this key information, ensuring that the chunk retains the core semantic content of the document. While chunking helps reduce the input length, it may also result in an uneven distribution of meaningful content across chunks. Thus, it is crucial to re-emphasize the importance of these keyphrases within the chunk to maintain semantic integrity. Our approach dynamically adjusts the representation of each chunk by assigning greater importance to keyphrase tokens, enabling the model to focus on the most relevant content during downstream processing. To achieve this, we label the tokens corresponding to the extracted keyphrases in the original text as keyphrase tokens $T_{k}$, while other tokens are labelled as non-keyphrase tokens $T_{nk}$. Then, we feed these chunked tokens $t$ into the embedding layer to obtain their embeddings. The chunk embedding $\boldsymbol{c}$ is then computed using a weighted average of these token embeddings, as defined in Formula 1:
$\begin{cases}w_{t}=\begin{cases}a,\text{t is $T_{k}$}\\ b,\text{t is $T_{nk}$}\end{cases}\\ \boldsymbol{c}=\frac{\sum{w_{t}*\boldsymbol{t}}}{\sum{w_{t}}}\end{cases}$ (1)
@@
Here, $w_{t}$ represents the weight assigned to each token $t$ in the chunk, where $a$ and $b$ are hyperparameters with $a>b.$ $\boldsymbol{t}$ denotes the embedding of token $t$, and $\boldsymbol{c}$ is the resulting chunk embedding that captures the weighted importance of keyphrase and non-keyphrase tokens. By assigning a higher weight $a$ to keyphrase tokens, we ensure that the resulting chunk representation emphasizes the most critical information while maintaining a compact input length. Finally, the chunk embeddings are fed into the Transformer-based model, allowing it to effectively leverage the enhanced chunk representations during long document classification or token-level classification tasks. This method not only preserves the semantic coherence of the document but also allows the model to retain meaningful context and relationships, ultimately enhancing its performance on long document tasks.